# Accelerating Transformer Pre-Training with 2:4 Sparsity

### v1

@@ Abstract
Training large transformers is slow, but recent innovations on GPU architecture give us an advantage. NVIDIA Ampere GPUs can execute a fine-grained 2:4 sparse matrix multiplication twice as fast as its dense equivalent. In the light of this property, we comprehensively investigate the feasibility of accelerating feed-forward networks (FFNs) of transformers in pre-training. First, we define a “flip rate” to monitor the stability of a 2:4 training process. Utilizing this metric, we propose three techniques to preserve accuracy: to modify the sparse-refined straight-through estimator by applying the masked decay term on gradients, to determine a feasible decay factor in warm-up stage, and to enhance the model’s quality by a dense fine-tuning procedure near the end of pre-training. Besides, we devise two techniques to practically accelerate training: to calculate transposable 2:4 masks by convolution, and to accelerate gated activation functions by reducing GPU L2 cache miss. Experiments show that our 2:4 sparse training algorithm achieves similar convergence to dense training algorithms on several transformer pre-training tasks, while actual acceleration can be observed on different shapes of transformer block apparently.
@@ 1 Introduction
  • To the best of our knowledge, this is the first report on end-to-end acceleration on pre-training transformers (Figure 6, Table 8). Experiments show that transformers pre-trained using our proposed sparse training scheme are comparable or even superior in accuracy to those trained with dense training methods (Table 4, 5).
@@ 2 Related Work
Sparsity for Inference Acceleration. Early methods include one-shot pruning ((Han et al., 2015, 2016; Lee et al., 2018; Mishra et al., 2021). Later methods (Evci et al. (2021); Zhou et al. (2021); Lasby et al. (2023)) suggest using dynamic sparse training (DST). Particularly, Zhou et al. (2021) proposes sparse-refined straight-through estimator (SR-STE) for 2:4 inference. Iterative magnitude-based pruning (IMP) methods (Chen et al., 2020, 2021; You et al., 2022), originated from the winning lottery ticket theory (Frankle & Carbin, 2019; Frankle et al., 2020), can also be viewed as a DST approach. All these methods only speedup the forward pass. They are insufficient to accelerate training.
2:4 Semi-Structured Sparsity for Training Acceleration. Accelerating training by 2:4 sparsity is hard, because both the forward and backward passes need to be accelerated. On some GPUs involving sparse tensor cores, 2:4-spMMs perform 2x faster than dense GEMMs (Mishra et al., 2021; BUSATO & POOL, ). In light of this, (Hubara et al., 2021) firstly proposes a transposable N:M mask to accelerate both output activations and input gradients computation in backward pass. Zhang et al. (2023) improve transposable mask to bi-directional mask (Bi-Mask) to further boost mask diversity. To accelerate calculating weight gradient via 2:4-spMM, an unbiased minimum-variance estimator (MVUE) is introduced (Chmiel et al., 2023). In addition, Xu et al. (2022) also achieve fully sparse training of CNNs using spatial similarity. However, all these works do not report end-to-end training speedups on 2:4 sparse tensor cores, and they are built for CNNs. Practical 2:4 training acceleration on transformers has not been reported so far.
Other Structured Sparsity for Training Acceleration. Structured sparsity means channel-wise pruning to dense networks. For instance, training a large model and then compressing it to be thinner or shallower seems effective (Li et al., 2020; Zhou et al., 2020), given a fixed accuracy requirement. However, it’s not memory-efficient due to the larger model’s redundancy. In addition, low-rank adaption proves to be an effective method to reduce fine-tuning costs (Hu et al., 2023), but it can’t accelerate the pre-training.
@@ 3.2 Fully Sparse Training with 2:4 Sparsity
GEMMs can be accelerated with structured sparsity methods. Particularly, 2:4 sparsity (Mishra et al., 2021) is a semi-structured sparsity pattern supported on NVIDIA Ampere architectures. A 2:4 sparse matrix partitions its elements into groups of four numbers, where each group has exactly two zeros. Depending on the direction of partition, there are row-wise 2:4 sparse matrix and column-wise 2:4 sparse matrix (Appendix A.1). With such sparsity, a GEMM $\boldsymbol{\mathbf{C}}=\boldsymbol{\mathbf{A}}\boldsymbol{\mathbf{B}}$ can be accelerated by 2x with the 2:4-spMM kernel if either $\boldsymbol{\mathbf{A}}$ is row-wise 2:4 sparse, or $\boldsymbol{\mathbf{B}}$ is column-wise 2:4 sparse (Appendix A.3).
To accelerate training, each GEMM in Eq. (1) should have one 2:4 sparse operand. In general, weights and output activation gradients are selected to be pruned due to relatively lower pruning-induced loss (Chmiel et al., 2023). That is,
@@
$\nabla_{\boldsymbol{\mathbf{X}}}=\boldsymbol{\mathbf{Z}}S_{w}(\boldsymbol{\mathbf{W}}),$ (3)
@@
In Eq. (2)-(4), $S_{wt},S_{w}$, and $S_{z}$ represent the pruning functions of $\boldsymbol{\mathbf{W}}^{\top},\boldsymbol{\mathbf{W}}$, and $\nabla_{\boldsymbol{\mathbf{Z}}}^{\top}$. They take a dense matrix as input, and outputs a 2:4 sparse matrix. By intuition, a pruning function should pick out the 2 elements with the max magnitudes in the adjoining 4 elements and zero out the rest. With hardware support, computing Eq. (2)-(4) can be theoretically 2x faster than computing Eq. (1). This method use 2:4-spMMs for all matrix multiplications in forward and backforward propagation, so we call it fully sparse training (FST). Note that Eq. (4) contains a straight-through estimator (STE). We will later explain it in Section 3.
@@
which they call as transposable masks (same as our defination in Section 5.1). In this manner, the backward pass share the same sparse weight matrix with the forward pass. In addition, the authors propose a 2-approximation method for generating such masks with claimed low computational complexity.
@@ 4 Accuracy Preserving Techniques
While the methods reviewed in Sec. 3 can successfully perform FST on small-scale models such as ResNets and DenseNets, it is not yet clear whether they can be directly applied to pre-train large transformers. It is challenging for FST to preserve the accuracy of dense training, since the weights and masks need to be learned jointly, which is a non-differentiable, combinatorial optimization problem. Moreover, unlike inference acceleration methods, FST has no pre-trained dense model to start with. In this section, we propose three practical techniques to improve the convergence of FST for transformers: transformer-specific masked decay and dense fine-tuning.
@@ 4.1 Flip Rate: Stability of Training
We measure flip rate change for dense training, STE and SR-STE with different $\lambda_{W}$ in Figure 1. For dense training, we compute the flip rate by pruning the dense weight in each iteration, despite the pruned weight is never used for training. In terms of flip rate, dense training is healthy: its $r_{t}$ exactly increases first before declines. If a training process consistently has higher flip rate than dense training, which we call as “flip rate explosion”, it may suffer from a loss in final accuracy due to unstable training. In practice, STE suffers from a flip rate explosion, while SR-STE takes effect by “freezing” masks of weights: by adding a decay term, it decrease the number of flips. This inhibition effect is related to the decay factor of SR-STE: the larger $\lambda_{W}$ is, the stronger the inhibition of flips is, and the smaller flip rate goes.
@@ 4.2 Transformer-Specific Masked Decay
Unlike Eq. (8) which imposes regularization directly on weights, we propose to add masked decay on gradients, i.e.,
@@
On SGD, applying the decay term on weights and on gradients are equivalent, while on popular optimizers like Adam and AdamW they are not. Specifically, Adam updates weights by
@@
where $\boldsymbol{\mathbf{u}}$ and $\boldsymbol{\mathbf{v}}$ are the first and second order momentum of $\boldsymbol{\mathbf{w}}$. Compared to Eq. (8), the masked decay regularization term in Eq. (10) would be later normalized by $\sqrt{\hat{\boldsymbol{\mathbf{v}}}_{t}}+\epsilon$ in Eq. (11), before it is subtracted from weights. In this way, each dimension receives a different intensity of decay (“masked decay”). More specifically, weights with larger gradients get smaller decay intensity, and vice versa.
@@
In FST, we periodically prune weights by their magnitudes. STE may cause the network to fall into such “dilemma points”, where a portion of pruned weights and unpruned weights have nearly the same total L1 norm. Thus, the network consistently oscillate between two possible masks $\boldsymbol{\mathbf{m}}_{1}$ and $\boldsymbol{\mathbf{m}}_{2}$, and it is unlikely to jump out the dilemma point itself. This process can be observed by the flip rate explosion of STE. On these occasions, we argue that an evenly masked decay applied on weights is insufficient to save the training from such “traps”. The weights don’t differentiate themselves after an update, so masks may oscillate back. By normalizing the weight gradients with $\sqrt{\hat{\boldsymbol{\mathbf{v}}}_{t}}+\epsilon$, our masked decay amplifies the regularization strength for the dimension with smaller gradient, pushing it towards zero. Then, the regularized dimension can no longer compete with other dimensions. So we effectively break the tie and push the training process out of the trap, towards a “healthier” state.
The comparison results between our masked decay defined in Eq. (10) and the conventional counterpart in Eq. (8) are shown in Figure 2. Results show that applying masked decay on weights takes no effect to inhibit flip rate explosion of STE while applying on gradients works well enough.
4.3 Fast Decay Factor Determination
The determination of the decay factor $\lambda_{W}$ in Eq. (10) is non-trivial: if $\lambda_{W}$ is excessively large, then the “peak” of the flip rate curve is not high enough; if $\lambda_{W}$ is too small, the “tail” of the curve is not low enough. Both do not provide a healthy training process. Besides, we find that $\lambda_{W}$ values for CNNs and other small-scale networks differ significantly from those for transformers, while on transformers, optimal $\lambda_{W}$ can span up to three orders of magnitude (Table 1).
@@ 4.3 Fast Decay Factor Determination
    Grid search on the warm-up stage of training. That is, for each $\lambda_{W}$ value in a candidate set, sample a corresponding flip rate of the sparse network from a small number of training steps. Note that sampling in early training stage is enough to obtain a representative flip rate specific to a sparse network.
@@ 4.4 Dense Fine-Tuning
Some recent works like STEP (Lu et al., 2023) utilize dense pre-training rather than dense fine-tuning, which means a dense network is initially trained for a period of time before being switched to a sparse one. However, we argue that dense pre-training is meaningless in our FST process. As described in Section 4.1, the peak of the flip rate curve should be sufficiently high to explore connection modes, so what matters most to the flip rate is the magnitudes of weights, which are the key to determine if connections are built or demolished. In this regard, both FST and dense pre-training are capable of delivering proper gradient magnitudes, so dense pre-training is a waste. The precise gradients are generally more necessary in the later stages of training, where the flip rate of the dense network comes to its tail. Figure 3 visualizes the loss curve of pre-training BERT-base, where dense pre-train obtains nearly the same result as the naive SR-STE method. From this, we propose the following insight:
@@ 5 Training Acceleration Techniques
For transformers, the forward pass of our FST involves pruning weights in feed-forward networks with transposable 2:4 masks and then performing normal forward propagation. During the backward propagation in our FST, the gradients of input activations and weight gradients in FFNs are derived by Eq. (3) and (4), respectively. Note that we also utilize MVUE to prune gradients of output activations, i.e., Eq. (6). Compared to dense training, our FST replaces all the GEMMs in FFNs with 2:4-spMMs that theoretically perform 2x faster than their dense counterparts on GPUs within sparse tensor cores.
@@
Below, we show our methods to accelerate these operations.

### v3

@@ Abstract
Training large transformers is slow, but recent innovations on GPU architecture give us an advantage. NVIDIA Ampere GPUs can execute a fine-grained 2:4 sparse matrix multiplication twice as fast as its dense equivalent. In the light of this property, we comprehensively investigate the feasibility of accelerating feed-forward networks (FFNs) of transformers in pre-training. First, we define a “flip rate” to monitor the stability of a 2:4 training process. Utilizing this metric, we propose three techniques to preserve accuracy: to modify the sparse-refined straight-through estimator by applying the masked decay term on gradients, to determine a feasible decay factor in warm-up stage, and to enhance the model’s quality by a dense fine-tuning procedure near the end of pre-training. Besides, we devise two techniques to practically accelerate training: to calculate transposable 2:4 masks by convolution, and to accelerate gated activation functions by reducing GPU L2 cache miss. Experiments show that our 2:4 sparse training algorithm achieves similar convergence to dense training algorithms on several transformer pre-training tasks, while actual acceleration can be observed on different shapes of transformer block apparently. Our toolkit is available at https://github.com/huyz2023/2by4-pretrain.
@@ 1 Introduction
  • $\bullet$
    To the best of our knowledge, this is the first report on end-to-end acceleration on pre-training transformers (Figure 7, Table 11). Experiments show that transformers pre-trained using our proposed sparse training scheme are comparable or even superior in accuracy to those trained with dense training methods (Table 5, 6).
@@ 2 Related Work
Sparsity for Inference Acceleration
Early methods include one-shot pruning (Han et al., 2015, 2016; Lee et al., 2018; Mishra et al., 2021). Later methods (Evci et al., 2021; Zhou et al., 2021; Lasby et al., 2023) suggest using dynamic sparse training (DST). Particularly, Zhou et al. (2021) proposes sparse-refined straight-through estimator (SR-STE) for 2:4 inference. Iterative magnitude-based pruning (IMP) methods (Chen et al., 2020, 2021; You et al., 2022), originated from the winning lottery ticket theory (Frankle & Carbin, 2019; Frankle et al., 2020), can also be viewed as a DST approach. All these methods only speedup the forward pass. They are insufficient to accelerate training.
2:4 Semi-Structured Sparsity for Training Acceleration
Accelerating training by 2:4 sparsity is hard, because both the forward and backward passes need to be accelerated. On some GPUs involving sparse tensor cores, 2:4-spMMs perform 2x faster than dense GEMMs (Mishra et al., 2021; BUSATO & POOL, ). In light of this, (Hubara et al., 2021) firstly proposes a transposable N:M mask to accelerate both output activations and input gradients computation in backward pass. Zhang et al. (2023) improve transposable mask to bi-directional mask (Bi-Mask) to further boost mask diversity. To accelerate calculating weight gradient via 2:4-spMM, an unbiased minimum-variance estimator (MVUE) is introduced (Chmiel et al., 2023). In addition, Xu et al. (2022) also achieve fully sparse training of CNNs using spatial similarity. However, all these works do not report end-to-end training speedups on 2:4 sparse tensor cores, and they are built for CNNs. Practical 2:4 training acceleration on transformers has not been reported so far.
Other Structured Sparsity for Training Acceleration
Structured sparsity means channel-wise pruning to dense networks. For instance, training a large model and then compressing it to be thinner or shallower seems effective (Li et al., 2020; Zhou et al., 2020), given a fixed accuracy requirement. However, it’s not memory-efficient due to the larger model’s redundancy. In addition, low-rank adaption proves to be an effective method to reduce fine-tuning costs (Hu et al., 2023), but it can’t accelerate the pre-training.
@@ 3.2 Fully Sparse Training with 2:4 Sparsity
GEMMs can be accelerated with structured sparsity. Particularly, 2:4 sparsity (Mishra et al., 2021) is a semi-structured sparsity pattern supported on NVIDIA Ampere architectures. A 2:4 sparse matrix partitions its elements into groups of four numbers, where each group has exactly two zeros. Depending on the direction of partition, there are row-wise 2:4 sparse matrix and column-wise 2:4 sparse matrix; see Section A.1. With such sparsity, a GEMM $\boldsymbol{\mathbf{C}}=\boldsymbol{\mathbf{A}}\boldsymbol{\mathbf{B}}$ can be accelerated by 2x with the 2:4-spMM kernel if either $\boldsymbol{\mathbf{A}}$ is row-wise 2:4 sparse, or $\boldsymbol{\mathbf{B}}$ is column-wise 2:4 sparse.
To accelerate training, each GEMM in Equation 1 should have one 2:4 sparse operand. In general, weights and output activation gradients are selected to be pruned due to relatively lower pruning-induced loss (Chmiel et al., 2023). That is,
@@
$\nabla_{\boldsymbol{\mathbf{X}}}=\nabla_{\boldsymbol{\mathbf{Z}}}S_{w}(\boldsymbol{\mathbf{W}}),$ (3)
@@
In Equations 2, 3 and 4, $S_{wt},S_{w}$, and $S_{z}$ represent the pruning functions of $\boldsymbol{\mathbf{W}}^{\top},\boldsymbol{\mathbf{W}}$, and $\nabla_{\boldsymbol{\mathbf{Z}}}^{\top}$. They take dense matrices as input, and outputs 2:4 sparse matrices. By intuition, a pruning function picks out the 2 elements with the max magnitudes in the adjoining 4 elements and zero out the rest. With hardware support, computing Equations 2, 3 and 4 can be theoretically 2x faster than Equation 1. This method use 2:4-spMMs for all matrix multiplications in forward and backward propagation, so we call it fully sparse training (FST). Note that Equation 4 contains a straight-through estimator (STE), which we will explain later.
@@
which they call as transposable masks (same as our defination in Section 5.1). In this manner, the backward pass share the same sparse weight matrix with the forward pass. The authors also propose a 2-approximation method for generating such masks with claimed low computational complexity.
@@ 4 Accuracy Preserving Techniques
While the methods reviewed in Section 3 can successfully perform FST on small-scale models such as ResNet and DenseNet, it is not clear whether they can be directly applied to pre-train large transformers. It is challenging for FST to preserve the accuracy of dense training, since the weights and masks need to be learned jointly, which is a non-differentiable, combinatorial optimization problem. Moreover, unlike inference acceleration methods, FST has no pre-trained dense model to start with. In this section, we propose three practical techniques to improve the convergence of FST for transformers: transformer-specific masked decay, Fast decay factor determination and dense fine-tuning.
@@ 4.1 Flip Rate: Stability of Training
Table 1: Training results of different $\lambda_{W}$ on Transformer-base. As $\lambda_{W}$ increases from 0 to 2e-4, accuracy first rises and then drops, which means that $\lambda_{W}$ should be neither too big nor too small to reach the optimal results.
$\lambda_{W}$ Avg epoch loss Val loss Test BLEU
Dense 4.558 3.978 26.15
0 (STE) 4.76 4.164 24.98
6e-7 4.684 4.079 25.68
6e-6 4.626 4.033 25.81
2e-6 4.64 4.041 25.94
2e-5 4.642 4.049 25.74
2e-4 4.662 4.06 25.62
@@
We measure flip rate change for dense training, STE and SR-STE with different $\lambda_{W}$ in Figure 1. For dense training, we compute the flip rate by pruning the dense weight in each iteration, despite the pruned weight is never used for training. In terms of flip rate, dense training is healthy: its $r_{t}$ exactly increases first before declines. If a training process consistently has higher flip rate than dense training, which we call as “flip rate explosion”, it may suffer from a loss in final accuracy due to unstable training; see Table 1. In practice, STE suffers from a flip rate explosion, while SR-STE takes effect by “freezing” masks of weights: by adding a decay term, it decrease the number of flips. This inhibition effect is related to the decay factor of SR-STE: the larger $\lambda_{W}$ is, the stronger the inhibition of flips is, and the smaller flip rate goes.
@@ 4.2 Transformer-Specific Masked Decay
Unlike Equation 8 which imposes regularization directly on weights, we propose to add masked decay on gradients, i.e.,
@@
On SGD, applying decay on weights and on gradients are equivalent, but on popular optimizers like Adam and AdamW they aren’t. Specifically, Adam updates weights by
@@
where $\boldsymbol{\mathbf{u}}$ and $\boldsymbol{\mathbf{v}}$ are the first and second order momentum of $\boldsymbol{\mathbf{w}}$. Compared to Equation 8, the masked decay regularization term in Equation 10 would be later normalized by $\sqrt{\hat{\boldsymbol{\mathbf{v}}}_{t}}+\epsilon$ in Equation 11, before it is subtracted from weights. In this way, each dimension receives a different intensity of decay (“masked decay”). More specifically, weights with larger gradients get smaller decay intensity, and vice versa.
In FST, we periodically prune weights by their magnitudes. STE may cause the network to fall into such “dilemma points”, where a portion of pruned weights and unpruned weights have nearly the same L1 norm. Thus, the network consistently oscillate between two possible masks $\boldsymbol{\mathbf{m}}_{1}$ and $\boldsymbol{\mathbf{m}}_{2}$, and is unlikely to jump out the dilemma itself. To illustrate this, we split each weight matrix by small $4\times 4$ blocks. We count each block’s cumulative flip number and measure the ”L1 norm gap” by $g_{i}=\left\lVert\boldsymbol{\mathbf{m}}_{1}\odot\boldsymbol{\mathbf{w}}_{i}\right\rVert_{1}-\left\lVert\boldsymbol{\mathbf{m}}_{2}\odot\boldsymbol{\mathbf{w}}_{i}\right\rVert_{1}$, where $\boldsymbol{\mathbf{w}}_{i}$ is the $i$-th $4\times 4$ weights, $\boldsymbol{\mathbf{m}}_{1}\odot\boldsymbol{\mathbf{w}}_{i}$ and $\boldsymbol{\mathbf{m}}_{2}\odot\boldsymbol{\mathbf{w}}_{i}$ have the first and second largest L1-norm among different pruning binary masks. The selected mask is most likely to oscillate between $\mathbf{m}_{1}$ and $\mathbf{m}_{2}$, especially when $g_{i}$ is small. In STE, there exists more $4\times 4$ blocks with high flip num and low ”L1 norm gap”; see Figure 2. This results in overall flip rate explosion of STE.
Figure 2: Scatter plots of cumulative flip number and L1 norm gap $g_{i}$ on every $4\times 4$ block. All results are selected on Transformer-base, with epoch=20. (a) shows the result of dense model. (b)-(d) shows that of masked decaying on gradients, no decaying, and masked decaying on weights. Also, we do it on purpose to choose an extremely large $\lambda_{W}$ for SR-STE.
@@
On these occasions, we argue that an evenly masked decay applied on weights is insufficient to save the training from such “traps”. The weights don’t differentiate themselves after an update, so masks may oscillate back. By normalizing the weight gradients with $\sqrt{\hat{\boldsymbol{\mathbf{v}}}_{t}}+\epsilon$, our masked decay amplifies the regularization strength for the dimension with smaller gradient, pushing it towards zero. Then, the regularized dimension can no longer compete with other dimensions. So we effectively break the tie and push the training process out of the trap, towards a “healthier” state.
The comparison results between our masked decay defined in Equation 10 and the conventional counterpart in Equation 8 are shown in Figure 3. Results show that applying masked decay on weights takes no effect to inhibit flip rate explosion of STE, while applying on gradients works fine.
4.3 Fast Decay Factor Determination
The determination of the decay factor $\lambda_{W}$ in Equation 10 is non-trivial: if $\lambda_{W}$ is excessively large, then the “peak” of the flip rate curve is not high enough; if $\lambda_{W}$ is too small, the “tail” of the curve is not low enough. Both do not provide a healthy training process. Besides, we find that $\lambda_{W}$ values for CNNs and other small-scale networks differ significantly from those for transformers, while on transformers, optimal $\lambda_{W}$ can span up to three orders of magnitude (Table 2).
@@ 4.3 Fast Decay Factor Determination
    Grid search on the warm-up stage of training. For each $\lambda_{W}$ value in a candidate set, sample a corresponding flip rate of the sparse network from a small number of training steps. Note that sampling in early training stage is enough to obtain a representative flip rate specific to a sparse network.
@@ 4.4 Dense Fine-Tuning
While previous work (Han et al., 2017) suggest to switch between sparse and dense training stages, some recent works like STEP (Lu et al., 2023) utilize dense pre-training rather than dense fine-tuning, which means a dense network is initially trained for a period of time before being switched to a sparse one. However, we argue that dense pre-training is meaningless in our FST process. As described in Section 4.1, the peak of the flip rate curve should be sufficiently high to explore connection modes, so what matters most to the flip rate is the magnitudes of weights, which are the key to determine if connections are built or demolished. In this regard, both FST and dense pre-training are capable of delivering proper gradient magnitudes, so dense pre-training is a waste. The precise gradients are generally more necessary in the later stages of training, where the flip rate of the dense network comes to its tail. Figure 4 visualizes the loss curve of pre-training BERT-base, where dense pre-train obtains nearly the same result as the naive SR-STE method. From this, we propose the following insight:
@@ 5 Training Acceleration Techniques
For transformers, the forward pass of FST involves pruning weights in FFNs with transposable 2:4 masks and then performing normal forward propagation. During backward propagation in FST, the gradients of input activations and weight gradients in FFNs are derived by Equation 3 and (4), respectively. Note that we also utilize MVUE to prune gradients of output activations, i.e., Equation 6. Compared to dense training, our FST replaces all the GEMMs in FFNs with 2:4-spMMs that theoretically perform 2x faster than their dense counterparts on GPUs within sparse tensor cores.
@@
Below, we show our methods to accelerate these operations, the main workflow of which is shown in Appendix B.